# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v1

@@ 3.1 Task
This task was introduced in Mamba gu2024mambalineartimesequencemodeling to ensure that models are able to ignore distractors, which transformers with static positional encoding like RoPE su2023roformerenhancedtransformerrotary are unable to do. Intuitively, the model has to "compress" the distances between tokens, and increment its position within the cognitive map (here the number line) every time it sees a non blank token. Solving this task requires a flexible gating mechanism that baseline transformers do not possess. The model receives a sequence of symbols and must copy it entirely while ignoring a blank token $B:CFEABABBBCBF\to CFEAACF$.
@@
3.2 Baseline models
In all our experiments, we will compare our models to RoPE su2023roformerenhancedtransformerrotary and CoPE golovneva2024contextualpositionencodinglearning. RoPE is used as our baseline Transformer since it is the most popular (fixed) position encoding architecture, therefore incapable of learning input-dependent matrices, and as we will see, structure-based problems. CoPE is an enhanced Transformer architecture that learns a gating mechanism akin to Mamba SSMs, allowing them to adapt their position encoding by compressing relative distances between tokens, and solve structure dependent tasks such as selective copy. CoPE will serve as a baseline for structure learning models.
@@ 3.5.1 MAmPa as a Model of Episodic Memory (EM) and Working Memory (WM)
In Whittington2025, the authors introduce two related models of episodic (EM) and working memory (WM) that both learn cognitive maps using RNNs with input-dependent matrices. Discrete SSMs can be regarded as RNNs with no activation function, and since selective SSMs learn input-dependent matrices, EM and WM models can naturally be regarded as selective SSMs. The EM-RNN model behaves exactly like TEM Whittington770495, factorizing structure and content in two distinct neural populations, but only updates its position with an input-dependent RNN $p_{t}:=W_{a_{t}}p_{t-1}$. It then queries a Hopfield memory module ramsauer2021hopfieldnetworksneed with its updated position to retrieve passed observations. The RNN in TEM can be seen as an SSM with two subtle differences: (1) it uses no external input, i.e. $\forall t,x_{t}=0$; (2) a learnable embedding $p_{*}$ representing the cognitive map’s zero coordinate must be introduced to avoid the state being always null:
$p_{t}$ $=\exp(\Delta_{t}A)p_{t-1}=\exp\left(\omega S\sum_{0}^{t}\Delta_{s}\right)p_{*}$ (10)
@@
where $C$ is an optional learnable parameter to project the updated position out of the cognitive map like in standard SSMs, but setting $C:=\mathbf{1}_{n}$ implies that $q_{t}=p_{t}$, which is analogous to TEM/EM RNNs (fig. 1(c)). Moreover, compared to RNN based models, this formulation allows to compute path integration directly in the Lie algebra ($A\sum_{0}^{t}\Delta_{s}$) and avoid sequential matrix multiplication.
WM-RNNs directly updates the internal state with input-dependent matrices, which is exactly what selective SSMs do. As such, current selective SSMs with the appropriate matrix structure (skew-symmetric in the case of navigation, eq. 3) naturally model working memory (fig. 1(d)). This allows us to unify SSMs with EM and WM models, where the SSM’s inner state-space can now be understood as a cognitive map.
In the previous section, analyzing SSMs through Lie-group theory has allowed us to understand their current limitations, and that one only needs to switch to block-diagonal matrices to theoretically be able to learn a cognitive map. However, people decided to keep diagonal matrices because it dramatically reduces the computational load. By treating $A$ as a vector, one can avoid computing a matrix exponential and also dramatically reduce the memory footprint of a parallel scan, by computing vector-vector multiplications instead of matrix-vector’s. Even then, Mamba uses a very small inner state $N\approx 16$, showing how difficult it is to scale these architectures. In Mamba2 dao2024transformersssmsgeneralizedmodels, the authors decide to reduce $A:=a\in\mathbb{R}$ in order to scale the inner state’s size $N=64$ to modern Transformer standards. This highlights a fatal limitations of parallel-selective SSMs: they can be parallelized at the cost of loosing any hope of learning a cognitive map (see sec. 10.1 for further details). This is the reason why we decided to move away from SSMs and adapt Transformers to learn cognitive maps, as models of episodic (MapFormer-EM) and working memory (MapFormer-WM).
@@ 3.7 Relationship between MapFormers and MAmPa: Linear MapFormers are selective SSMs
If we remove the matrices $A_{t}$, the matrix $M:=(m_{ij}):m_{ij}=C_{j}^{T}B_{i}:=q_{j}^{T}k_{i}$ can be seen as the key-query similarity matrix $QK^{T}$ without softmax normalization, as in linear transformers. Equation 16 is fundamental as it characterizes exactly how actionable representations can be implemented in transformers, where matrix $A_{i\to j}^{\times}=A_{j}...A_{i+1}$ represents the model’s path integration between tokens $i$ and $j$.
@@ 5.1 Selective copy
Table 1: Selective copy accuracy. IID sequences are generated with $128$ tokens to copy and $128$ blank tokens $B$ to ignore. OOD-d (dense) and OOD-s (sparse) vary the amount of blank tokens to ignore to $64$ and $256$, respectively. RoPE solves the task IID but fails to generalize OOD, while as expected, both CoPE and MapFormer can solve it, showing that the latter can also learn gating mechanisms. MapEM-o and MapEM-s cannot solve this task, since it requires both "sensory" inputs (the tokens to copy) and structure (the tokens to ignore).
In line with golovneva2024contextualpositionencodinglearning, we find that RoPE perfectly learns in distribution but completely fails to generalize, while as expected, both CoPE and our models manage to solve the task on OOD samples, since moving forward on a finite segment or a circle are equivalent. Because this task involves copying relevant tokens, it cannot be solved without sensory inputs (the token to copy), and therefore MapEM-s, our model relying on structure alone, cannot solve it.
@@ 5.2 Forced Grid Navigation
Table 2: 1D-2D grid navigation accuracy. RoPE does not solve the task since it does not learn input-dependent matrices. CoPE performs better but does not solve the task, even in 1D, showing that gating is not enough to learn 1D cognitive maps, since it does not allow to learn the inverse of an action (forward/backward). Even with up to 4 layers, both RoPE and CoPE fail to solve the task, showing that without the appropriate structure, deeper models cannot learn a cognitive map. All MapFormers model solved the task and perfectly generalize OOD, even if performances slightly degrade in 2D on longer sequences (OOD-s), since the amount of coordinates to represent increases quadratically. As expected MapEM-o, a transformer with no positional embeddings, performs even worse that RoPE, since it does not encode position.
@@
We first tested our models on a 1D navigation task, which has a subtle difference with gating. Even though the underlying structure is 1-dimensional, the model must be able to move forward or backward – i.e. representing an action and its inverse – while gating only compresses the ordinal order and can only move forward. Therefore, we hypothesize that CoPE, which learns a gating structure, should not be able to solve this task. As expected, all MapFormer models solve the task, but both RoPE and CoPE fail (tab. 2), showing a key limitation of gating even in 1D.
@@
As expected, our three model variants perfectly solve the task in 2D and generalize it to shorter or longer sequences. Even though CoPE does significantly better than RoPE, it fails to reach a perfect accuracy on both IID and OOD samples. Note that because 2D positions require more neurons to represent the two orthogonal axis, performances degrade faster on longer sequences than in 1D (tab. 2).
For both RoPE and CoPE, increasing the number of layers does not change the model’s performance, even though CoPE seems to benefit from more layers in 1D, probably because it helps the model to learn a backward operation that cannot be approximated with a single layer. Nevertheless, these results suggest that structure cannot be approximated with deeper networks without the appropriate inductive bias.
@@
These results show the importance of learning input-dependent positional encoding in order to solve structure-dependent tasks with Transformers. CoPE, which only learns a gating mechanism, can solve the Selective Copy task but fails to learn a cognitive map, compared to MapFormers that perfectly generalize OOD in all tasks, demonstrating the expressive power of input-dependent matrices to solve structure-dependent problems.
@@ 5.3 EM is computationally more expensive but scales better than WM
To test the impact of model size and sequence length, we vary the head size from $16$ to $128$ and sequence length from $16$ to $384$, respectively, while keeping a small vocabulary size of $10$. To test the impact of vocabulary on recall, we varied the number of objects from 10 to 10000, and train small models with head size 32 on short sequences of size 16. Results are reported in fig. 4
5.4 Observations are vectors, actions are matrices
@@ 6 Discussion
We introduced MapFormers, a novel class of Transformer-based architectures designed to learn cognitive maps and perform parallel path integration in a fully self-supervised setup. Our approach addresses the key limitation of previous RNN-based models by leveraging the parallel processing ability of Transformers, which is essential for scaling to complex, long-sequence problems. We developed two variants, the MapFormer-WM (Working Memory) and MapFormer-EM (Episodic Memory), which unify relative and absolute positional embeddings, respectively, as computational models of these distinct memory systems. Critically, we extended the theoretical framework of actionable representations dorrell2023actionableneuralrepresentationsgrid, drawing on Lie-group theory, to enable this efficient, parallel path integration for large-scale structure learning.
Modern AI systems struggle with Out-of-Distribution (OOD) generalization and require orders of magnitude more data to achieve performance on par with humans. This limited OOD generalization can be explained by a misrepresentation of the true structure underlying the train and test data, suggesting that extracting and representing this structure in the form of a cognitive map offers a potential remedy. In our experiments, we found that MapFormers achieve almost perfect generalization to longer and sparser sequences, significantly outperforming baseline models. For problems that required only simple gating capabilities, MapFormers and models using CoPE golovneva2024contextualpositionencodinglearning achieved similar OOD generalization, as expected; however, for more complex problems which require the creation of a cognitive map, only MapFormers successfully generalized OOD. This demonstrates the important role of input-dependent positional encoding — endowed with the appropriate group structure — for structure-based OOD generalization, distinguishing it from generalization that might merely rely on sophisticated heuristics.
@@
MapFormers overcome the computational bottleneck of previous RNN-based EM/WM models Whittington2025; Whittington770495; whittington2022relatingtransformersmodelsneural, delivering efficiency and scalability comparable to standard Transformers. The main bottleneck of EM/WM RNNs is their slow processing time, but MapFormers require little extra compute compared to baseline Transformers. For problems that can be addressed with rotational operation, using $2\times 2$ rotation matrices allowed us to avoid computing a matrix exponential and matrix-vector products (eq. 18), while remaining compatible with FlashAttention for MapWM models. Our experiments show the effectiveness of MapFormers, trained on significantly longer sequences than those explored in previous work Whittington2025.
Compared to previous actionable models Whittington2025; Whittington770495; whittington2022relatingtransformersmodelsneural, using matrix exponential to learn input-dependent matrices allows us to analyze sequence processing models through Lie-group theory. This framework is useful to better characterize the type of structures that are learnable by the model. Furthermore, MapFormers provide greater flexibility compared to previous actionable models because they eliminate the need to assume the underlying structure. For instance, prior models required either the explicit knowledge of the number of possible actions or the irreducible matrix representations (the irreps) of the underlying group structure, which is often a prohibitive requirement in most applications.
The exponential map – needed to learn input-dependent matrices and therefore cognitive maps – arises in discrete SSMs via the integration of infinitesimal recurrent changes over a varying time interval. This explains how a model with fixed synaptic connections $A$ can learn a cognitive map by just varying the duration at which it exposes the input to neural transformations, giving rise to the matrix $\overline{A_{t}}:=\exp(\Delta_{t}A)$ (eq. 8), indicating that the neural mechanisms allowing animals to build cognitive maps is easier than we might think, since it naturally emerges in continuous, selective recurrent networks.
@@
The Transformer architecture is generic; it has been used in a variety of different tasks from language modeling to vision and protein folding Jumper2021; siméoni2025dinov3. The mechanism for learning a cognitive map, as our formalism shows, is also generic. The combination of the two in MapFormers suggests that non-spatial cognitive maps can thus be learned in a variety of different tasks. For example, in the case of language modeling, learning a cognitive map would mean disentangling the structure of sentences (syntax) from their content (semantics).
@@
While in this work we explored with only rotation matrices, for computational efficiency, more complex group structures (including non-commutative) can be learned by MapFormers. This might affect inference speed, making a tradeoff between expressibility and computing time (sec. 10.4).

### v2

@@ 3.1 Task
This task was introduced in Mamba gu2024mambalineartimesequencemodeling and later used in CoPE golovneva2024contextualpositionencodinglearning to ensure that models are able to ignore distractors, which transformers with static positional encoding like RoPE su2023roformerenhancedtransformerrotary are unable to do. The model receives a sequence of symbols and must copy it entirely while ignoring a blank token $B:CFEABABBBCBF\to CFEAACF$. Intuitively, the model has to "compress" the distances between tokens, and increment its position within the cognitive map (here the number line) every time it sees a non blank token (incrementing the number of tokens to copy by one). Solving this task requires a flexible gating mechanism in positional encoding that baseline transformers do not possess.
@@ 3.4.1 MAmPa as a Model of Episodic Memory (EM) and Working Memory (WM)
In Whittington2025, the authors introduce two related models of episodic (EM) and working memory (WM) that both learn cognitive maps using RNNs with input-dependent matrices. Here, we explain why selective, discrete State-Space Models (SSMs) can be viewed as EM and WM models. First, note that discrete SSMs can be regarded as RNNs with no activation function. Second, selective SSMs learn input-dependent matrices via an exponential map (eq. 8), meaning that their inner state captures the task’s latent structure, as in EM/WM models. Therefore, EM and WM models can naturally be regarded as selective SSMs.
SSM as a model of episodic memory
The EM-RNN model behaves exactly like TEM Whittington770495, factorizing structure and content in two distinct neural populations, but only updates its position with an input-dependent RNN $p_{t}:=W_{a_{t}}p_{t-1}$. It then queries a Hopfield memory module ramsauer2021hopfieldnetworksneed with its updated position to retrieve passed observations. The RNN in TEM can be seen as an SSM with subtle differences: (1) it uses no external input, i.e. $\forall t,x_{t}=0$; (2) a learnable embedding $p_{*}$ representing the cognitive map’s zero coordinate must be introduced to avoid the state being always null; (3) SSM’s inner state must be projected back to its final state ($y_{t}=Ch_{t}$, eq.8), which implies in this case that the final position $q_{t}:=Cp_{t}$ can be obtained with an optional learnable projection $C$; (4) as in TEM, the memory module is queried with the conjunction $g_{t}$ of what has been seen and where it was observed (fig. 1(c)):
$p_{t}$ $=\exp(\Delta_{t}A)p_{t-1}=\exp\left(\omega S\sum_{s=0}^{t}\Delta_{s}\right)p_{*}$ (10)
@@
Setting $C:=\mathbf{1}_{n}$ implies that $q_{t}=p_{t}$, which is analogous to TEM/EM RNNs. Moreover, compared to RNN based models, this formulation allows to compute path integration directly in the Lie algebra ($A\sum_{0}^{t}\Delta_{s}$) and avoid sequential matrix multiplication.
SSM as a model of working memory
WM-RNNs directly updates the internal state with input-dependent matrices, which is precisely what selective SSMs do. As such, current selective SSMs with the appropriate matrix structure (skew-symmetric in the case of navigation, eq. 3) naturally model working memory (fig. 1(d)).
Together, this allows us to view SSMs as both EM and WM models, where the inner state of an SSM can be interpreted as a cognitive map.
In the previous section, analyzing SSMs through Lie-group theory has allowed us to understand their current limitations, and that one only needs to switch to block-diagonal matrices to theoretically be able to learn a cognitive map. However, previous work decided to keep diagonal matrices because it dramatically reduces the computational load. By treating $A$ as a vector, one can avoid computing a matrix exponential and also dramatically reduce the memory footprint of a parallel scan, by computing vector-vector multiplications instead of matrix-vector’s. In Mamba2 dao2024transformersssmsgeneralizedmodels, it was further reduced to a scalar $A:=a\in\mathbb{R}$, which allows them to implement a fully parallelized version of selective SSMs but looses the ability to learn cognitive maps, as we have seen above. This highlights a fatal limitations of parallel-selective SSMs: they can be parallelized at the cost of loosing any hope of learning a cognitive map (see sec. 10.1 for further details). This is the reason why we decided to move away from SSMs and adapt Transformers to learn cognitive maps, as models of episodic (MapFormer-EM) and working memory (MapFormer-WM).
@@ 3.6 Relationship between MapFormers and MAmPa: Linear MapFormers are selective SSMs
If we remove the matrices $A_{t}$, the matrix $M:=(m_{ij});\;m_{ij}=C_{j}^{T}B_{i}:=q_{j}^{T}k_{i}$ can be seen as the key-query similarity matrix $QK^{T}$ without softmax normalization, as in linear transformers. Equation 16 is fundamental as it characterizes exactly how actionable representations can be implemented in transformers, where matrix $A_{i\to j}^{\times}=A_{j}...A_{i+1}$ represents the model’s path integration between tokens $i$ and $j$.
@@
3.7 Baseline models
In all our experiments, we will compare our MapFormer models to a classic Transformer architecture with static positional encoding, as well as two architectures learning input-dependent positional encoding serving as baselines for structure learning models.
Rotary Position Embedding (RoPE)
is used as our baseline Transformer since it is the most popular (fixed) position encoding architecture, and is analogous to our MapWM model without input selectivity, theoretically incapable of learning input-dependent matrices and structure-based problems su2023roformerenhancedtransformerrotary.
Contextualized Positional Encoding (CoPE)
is an enhanced Transformer architecture that updates its relative positions between queries and keys with a cumulative gating, akin to Mamba2 dao2024transformersssmsgeneralizedmodels. This allows CoPE to ignore irrelevant tokens and perfectly generalize OOD to several structure-dependent tasks. The cumulative gating of CoPE reminds path integration, but since the sigmoid gate can only take values between 0 and 1, the model cannot represent an action and its inverse, and should not be able to represent trajectories where the agent moves both forward and backward. Moreover, since CoPE only compresses relative distances, it should not be able to represent structure in higher dimensions, where temporal order does not matter golovneva2024contextualpositionencodinglearning.
Contextualized Equivariant Position Embedding (TAPE)
learns to modify its positional encoding after the attention mechanism, with input-dependent MLPs, resembling the input-dependent matrices of EM/WM actionable models Whittington2025. However, TAPE exhibits several issues: First, updating structure after the attention means that the latter cannot benefit from input-dependent positional encoding in the first layer, meaning that structure-based recall with a single-layer model is impossible. Second, updating values before updating structure means that values could be "polluted" by structurally irrelevant tokens, adding unwanted noise to the data. Third, TAPE’s positional embedding are updated based on the attention scores and not path integration, giving no theoretical guaranty to preserve structure over long ranges zhu2025rethinkingaddressinglanguagemodels.
@@ 5.1 Selective copy
Table 1: Selective copy accuracy. IID sequences are generated with $128$ tokens to copy and $128$ blank tokens $B$ to ignore. OOD-d (dense) and OOD-s (sparse) vary the amount of blank tokens to ignore to $64$ and $256$, respectively. RoPE solves the task IID but fails to generalize OOD, TAPE improves generalization on shorter sequences but not on long sequences. As expected, both CoPE and MapFormer can solve it, showing that the latter can also learn gating mechanisms. MapEM-o and MapEM-s cannot solve this task, since it requires both "sensory" inputs (the tokens to copy) and structure (the tokens to ignore).
In line with golovneva2024contextualpositionencodinglearning, we find that RoPE perfectly learns in distribution but completely fails to generalize. TAPE outperforms RoPE on the OOD-dense dataset, implying that it has learned a form of position interpolation that RoPE cannot learn, but its performances collapse on the OOD-sparse dataset, implying that TAPE did not learn to integrate actions over long ranges. As expected, both CoPE and MapFormers manage to solve the task on OOD samples, since moving forward on a finite segment or a circle are equivalent. Because this task involves copying relevant tokens, it cannot be solved without sensory inputs (the token to copy), and therefore MapEM-s, our model relying on structure alone, cannot solve it.
This experiment shows that all methods learning cumulative actions (i.e. path integration, CoPE and MapFormers) manage to perfectly generalize to longer and shorter sequences.
@@ 5.2 Forced Grid Navigation
Table 2: 1D-2D grid navigation accuracy. RoPE does not solve the task since it does not learn input-dependent matrices. CoPE performs better but does not solve the task, even in 1D, showing that gating is not enough to learn 1D cognitive maps, since it does not allow to learn the inverse of an action (forward/backward). Even with up to 4 layers, neither RoPE, TAPE nor CoPE manage to solve the task, showing that without the appropriate structural bias, deeper models cannot learn a cognitive map. All MapFormers model solved the task and perfectly generalize OOD, even if performances slightly degrade in 2D on longer sequences (OOD-s), since the amount of coordinates to represent increases quadratically. As expected MapEM-o, a transformer with no positional embeddings, performs even worse that RoPE, since it does not encode position.
@@
We first tested our models on a 1D navigation task, which has a subtle difference with gating. Even though the underlying structure is 1-dimensional, the model must be able to move forward or backward – i.e. representing an action and its inverse – while gating only compresses the ordinal order and can only move forward. Therefore, we hypothesize that CoPE, which learns a gating structure, should not be able to solve this task. As expected, all MapFormer models solve the task, but RoPE, CoPE and TAPE fail (tab. 2). The fact that CoPE fails in this task shows a key limitation of gating, even in 1D.
@@
As expected, our three MapFormer variants perfectly solve the task in 2D and generalize to shorter or longer sequences, unlike other architectures. Even though CoPE does significantly better than RoPE, it fails to reach a perfect accuracy on both IID and OOD samples. Note that because 2D positions require more neurons to represent the two orthogonal axis, performances of MapEM and MapWM degrade faster on longer sequences than in 1D (tab. 2).
@@
Increasing the number of layers for RoPE, CoPE, or TAPE does not change the model’s performance. CoPE seems to benefit from more layers in 1D only, probably because it helps the model to learn a backward operation that cannot be approximated with a single layer. TAPE’s performances improve IID with more layers, but increasing the model’s depth does not seem to impact OOD generalization. These results suggest that structure cannot be approximated with deeper networks without the appropriate inductive bias.
These experiments highlight the importance of learning input-dependent positional encoding in order to solve structure-dependent tasks with Transformers. CoPE, which only learns a gating mechanism, can solve the Selective Copy task but fails to learn a cognitive map, compared to MapFormers that perfectly generalize OOD in all tasks, demonstrating the expressive power of input-dependent matrices to solve structure-dependent problems.
@@ 5.3 EM is computationally more expensive but scales better than WM
To test the impact of model size and sequence length, we vary the head size from $16$ to $128$ and sequence length from $16$ to $384$, respectively, while keeping a small vocabulary size of $10$. To test the impact of vocabulary on recall, we varied the number of objects from 10 to 10000, and train small models with head size 32 on short sequences of size 16. Results are reported in fig. 4.
@@ 6 Discussion
We introduced MapFormers, a novel class of Transformer-based architectures designed to learn cognitive maps and perform parallel path integration in a fully self-supervised setup. Our approach addresses the key limitation of previous RNN-based models by leveraging the parallel processing ability of Transformers, which is essential for scaling to complex, long-sequence problems. We developed two variants, MapFormer-WM (Working Memory) and MapFormer-EM (Episodic Memory), which unify relative and absolute positional embeddings, respectively, as computational models of these distinct memory systems. Critically, we provide theoretical arguments from Lie-group theory – extending the theoretical framework of actionable representations dorrell2023actionableneuralrepresentationsgrid – to understand the class of structures that will be represented within a MapFormer’s positional encoding (sec. 3.2 &\& 10.4)
Modern AI systems struggle with out-of-distribution (OOD) generalization and require orders of magnitude more data to achieve performance on par with humans, which can be explained by a misrepresentation of the true structure underlying the train and test data. Previous work tried to solve this issue in Transformers by learning input-dependent relative positional encoding (TAPE, CoPE and PoPE zhu2025rethinkingaddressinglanguagemodels; golovneva2024contextualpositionencodinglearning; gopalakrishnan2025decouplingwhatwherepolar), achieving robust length generalization on some structure dependent tasks, but provide no tool to understand which type of structures can or cannot be learned by the model. Moreover, CoPE and PoPE only adjust the relative distances between tokens, making them unsuitable for 2-dimensional tasks or beyond. TAPE cannot learn structure with a single layer since it updates its positional encoding after the attention, nor perform path integration by aggregating the cumulative effect of actions. Compared to these methods, our work: provides theoretical tools to understand what structure can or cannot be learned by the model; allows to represent actions explicitly in a dedicated neural population (the Lie algebra); does not focus uniquely on relative positional encoding, highlighting a core difference between absolute and relative positional encoding on recall tasks (sec. 5.3).
In our experiments, we found that MapFormers achieved almost perfect generalization to longer and sparser sequences, significantly outperforming baseline models. For problems that only required simple gating capabilities (sec. 5.1), MapFormers and CoPE golovneva2024contextualpositionencodinglearning achieved similar OOD generalization, as expected by our theory. TAPE zhu2025rethinkingaddressinglanguagemodels, which does not explicitly aggregate the effect of actions via a cumulative sum – i.e. does not learn to path integrate – fails to generalize OOD on all tasks, even with more layers. This highlights the importance of aggregating local effects via a cumulative sum to coherently represent structure, independently of sequence length. For more complex problems which require the creation of a cognitive map (sec.5.2), only MapFormers successfully generalized OOD. This demonstrates the important role of updating positional encoding with input-dependent matrices – endowed with the appropriate group structure – for structure-based OOD generalization, distinguishing it from generalization that might merely rely on sophisticated heuristics.
@@
Compared to previous RNN-based EM/WM models Whittington2025; Whittington770495; whittington2022relatingtransformersmodelsneural, MapFormers exhibit several advantages. First, MapFormers overcome their computational bottleneck, requiring little extra compute compared to standard Transformers, maintaining their efficiency and scalability. Indeed, for problems that can be addressed with rotational operation (i.e., most cyclic abelian structures), using $2\times 2$ rotation matrices allowed us to avoid computing a matrix exponential and matrix-vector products (eq. 18), while remaining compatible with FlashAttention for MapWM models. Our experiments show the effectiveness of MapFormers, trained on significantly longer sequences than those explored in EM/WM RNNs so far. Furthermore, MapFormers provide greater flexibility compared to these models because they reduce the need to assume the underlying structure, and learn by themselves when to update or not the cognitive map. For instance, prior models Whittington2025; Whittington770495 required to explicitly know the number of possible actions and specify when to apply them, which is prohibitive in practice.
@@
Second, the scalability and structural flexibility makes MapFormers, like Transformers, generic architectures that could be used in a variety of different tasks from language modeling to vision and protein folding siméoni2025dinov3; Jumper2021. The mechanism for learning a cognitive map – as our formalism shows – is also generic, which suggests that non-spatial cognitive maps can thus be learned and studied in a variety of different tasks. For example, in the case of language modeling, learning a cognitive map would mean disentangling the structure of sentences (syntax) from their content (semantics).
@@
Third, while we mainly explored rotation matrices for computational efficiency, but more complex group structures can be learned by MapFormers. This allows notably to explore structures where the irreducible representations (irreps dorrell2023actionableneuralrepresentationsgrid, sec. 3.2) are unknown, like rotations in high dimensions or non-commutative groups. This might affect inference speed; that is, a tradeoff between expressibility and computing time would need to be found (sec. 10.4).
Finally, MapFormers provide a biologically-plausible interpretation of their dynamics. The exponential map – needed to learn input-dependent matrices and therefore cognitive maps – arises in Mamba SSMs gu2024mambalineartimesequencemodeling; dao2024transformersssmsgeneralizedmodels and MapFormers via the integration of infinitesimal recurrent changes over a varying time interval. This explains how a neural population with fixed synaptic connections $A$ can learn a cognitive map by just varying the duration at which it is exposed to the input ($\overline{A_{t}}:=\exp(\Delta_{t}A)$; eq. 8). This suggests a biologically plausible neural mechanisms that allows animals to build cognitive maps, which could naturally emerge in continuous, selective recurrent networks.
In summary, with MapFormers, we (1) address the issues of previous actionable models Whittington2025; Whittington770495; whittington2022relatingtransformersmodelsneural by enabling the learning of cognitive maps at scale; (2) unify absolute and relative positional encoding with models of Episodic Memory (EM) and Working Memory (WM); (3) give a comprehensive view of how to learn structure in Transformers, unlike previous attempts zhu2025rethinkingaddressinglanguagemodels; golovneva2024contextualpositionencodinglearning; gopalakrishnan2025decouplingwhatwherepolar; (4) empirically demonstrate that MapFormers are the only transformer-based architectures to perfectly generalize OOD on structure-dependent tasks.